# Grounding DINO: Marrying DINO with Grounded Pre-Training for Open-Set Object Detection

### v3

@@ Abstract
In this paper, we present an open-set object detector, called Grounding DINO, by marrying Transformer-based detector DINO with grounded pre-training, which can detect arbitrary objects with human inputs such as category names or referring expressions. The key solution of open-set object detection is introducing language to a closed-set detector for open-set concept generalization. To effectively fuse language and vision modalities, we conceptually divide a closed-set detector into three phases and propose a tight fusion solution, which includes a feature enhancer, a language-guided query selection, and a cross-modality decoder for cross-modality fusion. While previous works mainly evaluate open-set object detection on novel categories, we propose to also perform evaluations on referring expression comprehension for objects specified with attributes. Grounding DINO performs remarkably well on all three settings, including benchmarks on COCO, LVIS, ODinW, and RefCOCO/+/g. Grounding DINO achieves a $52.5$ AP on the COCO detection zero-shot transfer benchmark, i.e., without any training data from COCO. After fine-tuning with COCO data, Grounding DINO reaches $63.0$ AP. It sets a new record on the ODinW zero-shot benchmark with a mean $26.1$ AP. Code will be available at https://github.com/IDEA-Research/GroundingDINO.
Figure 1: (a) Closed-set object detection requires models to detect objects of pre-defined categories. (b) Previous work zero-shot transfer models to novel categories for model generalization. We propose to add Referring expression comprehension (REC) as another evaluation for model generalizations on novel objects with attributes. (c) We present an image editing application by combining Grounding DINO and Stable Diffusion [42]. Best view in colors.
00footnotetext: * This work was done when Shilong Liu, Feng Li, Hao Zhang, and Jie Yang were interns at IDEA. Corresponding authors.
@@ 1 Introduction
Understanding novel concepts is a fundamental capability of visual intelligence. In this work, we aim to develop a strong system to detect arbitrary objects specified by human language inputs, which we name as open-set object detection [^2] We view the terms open-set object detection, open-world object detection, and open-vocabulary object detection the same task in this paper. To avoid confusion, we always use open-set object detection in our paper.. The task has wide applications for its great potential as a generic object detector. For example, we can cooperate it with generative models for image editing (as shown in Fig. 1 (b)).
The key to open-set detection is introducing language for unseen object generalization [26, 1, 7]. For example, GLIP [26] reformulates object detection as a phrase grounding task and introduces contrastive training between object regions and language phrases. It shows a great flexibility for heterogeneous datasets and remarkable performance on both closed-set and open-set detection. Despite its impressive results, GLIP’s performance can be constrained since it is designed based on a traditional one-stage detector Dynamic Head [5]. As open-set and closed-set detection are closely related, we believe a stronger closed-set object detector can result in an even better open-set detector.
Motivated by the encouraging progress of Transformer-based detectors [58, 31, 24, 25], in this work, we propose to build a strong open-set detector based on DINO [58], which not only offers the state-of-the-art object detection performance, but also allows us to integrate multi-level text information into its algorithm by grounded pre-training. We name the model as Grounding DINO. Grounding DINO has several advantages over GLIP. First, its Transformer-based architecture is similar to language models, making it easier to process both image and language data. For example, as all the image and language branches are built with Transformers, we can easily fuse cross-modality features in its whole pipeline. Second, Transformer-based detectors have demonstrated a superior capability of leveraging large-scale datasets. Lastly, as a DETR-like model, DINO can be optimized end-to-end without using any hard-crafted modules such as NMS (Non-Maximum Suppression), which greatly simplifies the overall grounding model design.
Figure 2: Existing approaches to extending closed-set detectors to open-set scenarios. Note that some closed-set detectors can have only partial phases of the figure.
@@
Most existing open-set detectors are developed by extending closed-set detectors to open-set scenarios with language information. As shown in Fig. 2, a closed-set detector typically has three important modules, a backbone for feature extraction, a neck for feature enhancement, and a head for region refinement (or box prediction). A closed-set detector can be generalized to detect novel objects by learning language-aware region embeddings so that each region can be classified into novel categories in a language-aware semantic space. The key to achieving this goal is using contrastive loss between region outputs and language features at the neck and/or head outputs. To help a model align cross-modality information, some work tried to fuse features before the final loss stage. Fig. 2 shows that feature fusion can be performed in three phases: neck (phase A), query initialization (phase B), and head (phase C). For example, GLIP [26] performs early fusion in the neck module (phase A), and OV-DETR [56] uses language-aware queries as head inputs (phase B).
@@
We argue that more feature fusion in the pipeline enables the model to perform better. It is worth noting that retrieval tasks prefer a CLIP-like two-tower architecture which only performs multi-modality feature comparison at the end for efficiency. However, for open-set detection, the model is normally given both an image and a text input that specifies the target object categories or a specific object. In such a case, a tight (and early) fusion model is more preferred for a better performance [1, 26] as both image and text are available at beginning. Although conceptually simple, it is hard for previous work to perform feature fusion in all three phases. The design of classical detectors like Faster RCNN makes it hard to interact with language information in most blocks. Unlike classical detectors, the Transformer-based detector DINO has a consistent structure with language blocks. The layer-by-layer design enables it to interact with language information easily. Under this principle, we design three feature fusion approaches in the neck, query initialization, and head phases. More specifically, we design a feature enhancer by stacking self-attention, text-to-image cross-attention, and image-to-text cross-attention as the neck module. We then develop a language-guided query selection method to initialize queries for head. We also design a cross-modality decoder for the head phase with image and text cross-attention layers to boost query representations. The three fusion phases effectively help the model achieve better performance on existing benchmarks, which will be shown in Sec. 4.4.
Although significant improvements have been achieved in multi-modal learning, most existing open-set detection work evaluates their models on objects of novel categories, as shown in the left column of Fig. 1 (b). We argue that another important scenario, where objects are described with attributes, should also be considered. In the literature, the task is named Referring Expression Comprehension (REC) [34, 30] [^3] We use the term Referring Expression Comprehension (REC) and Referring (Object) Detection exchangeable in this paper.. We present some examples of REC in the right column of Fig. 1 (b). It is a closely related field but tends to be overlooked in previous open-set detection work. In this work, we extend open-set detection to support REC and also evaluate its performance on REC datasets.
@@
We conduct experiments on all three settings, including closed-set detection, open-set detection, and referring object detection, to comprehensively evaluate open-set detection performance. Grounding DINO outperforms competitors by a large margin. For example, Grounding DINO reaches a $52.5$ AP on COCO minival without any COCO training data. It also establishes a new state of the art on the ODinW [23] zero-shot benchmark with a $26.1$ mean AP.
The contributions of this paper are summarized as follows:
  1. 1.
    We propose Grounding DINO, which extends a closed-set detector DINO by performing vision-language modality fusion at multiple phases, including a feature enhancer, a language-guided query selection module, and a cross-modality decoder. Such a deep fusion strategy effectively improves open-set object detection.
  2. 2.
    We propose to extend the evaluation of open-set object detection to REC datasets. It helps evaluate the performance of the model with freeform text inputs.
  3. 3.
    The experiments on COCO, LVIS, ODinW, and RefCOCO/+/g datasets demonstrate the effectiveness of Grounding DINO on open-set object detection tasks.
@@
Base Detector Fusion Phases (Fig. 2) use CLIP Represent. Level (Sec. 3.4) COCO COCO LVIS ODinW RefCOCO/+/g
ViLD [13] Mask R-CNN [15] - sentence partial label partial label
RegionCLIP [62] Faster RCNN [39] - sentence partial label partial label
FindIt [21] Faster RCNN [39] A sentence partial label fine-tune
MDETR [18] DETR [2] A,C word fine-tune zero-shot fine-tune
DQ-DETR [46] DETR [2] A,C word zero-shot fine-tune
GLIP [26] DyHead [5] A word zero-shot zero-shot zero-shot
GLIPv2 [59] DyHead [5] A word zero-shot zero-shot zero-shot
OV-DETR [56] Deformable DETR [64] B sentence partial label partial label
@@
DetCLIP [53] ATSS [60] - sentence zero-shot zero-shot
OmDet [61] Sparse R-CNN [47] C sentence zero-shot
Grounding DINO (Ours) DINO [58] A,B,C sub-sentence zero-shot zero-shot zero-shot zero-shot
Table 1: A comparison of previous open-set object detectors. Our summarization is based on the experiments in their paper, but not the ability to extend their models to other tasks. It is worth noting that some related works may not (only) be designed for the open-set object detection initially, like MDETR [18] and GLIPv2[59], but we list them here for a comprehensive comparison with existing work. We use the term “partial label” for the settings, where models are trained on partial data (e.g. base categories) and evaluated on other cases.
@@ 2 Related Work
Detection Transformers. Grounding DINO is built upon the DETR-like model DINO [58], which is an end-to-end Transformer-based detector. DETR was first proposed in [2] and then has been improved from many directions [64, 33, 12, 5, 50, 17, 4] in the past few years. DAB-DETR [31] introduces anchor boxes as DETR queries for more accurate box prediction. DN-DETR [24] proposes a query denoising approach to stabilizing the bipartite matching. DINO [58] further develops several techniques including contrastive de-noising and set a new record on the COCO object detection benchmark. However, such detectors mainly focus on closed-set detection and are difficult to generalize to novel classes because of the limited pre-defined categories.
Open-Set Object Detection. Open-set object detection is trained using existing bounding box annotations and aims at detecting arbitrary classes with the help of language generalization. OV-DETR [57] uses image and text embedding encoded by a CLIP model as queries to decode the category-specified boxes in the DETR framework [2]. ViLD [13] distills knowledge from a CLIP teacher model into a R-CNN-like detector so that the learned region embeddings contain the semantics of language. GLIP [11] formulates object detection as a grounding problem and leverages additional grounding data to help learn aligned semantics at phrase and region levels. It shows that such a formulation can even achieve stronger performance on fully-supervised detection benchmarks. DetCLIP [53] involves large-scale image captioning datasets and uses the generated pseudo labels to expand the knowledge database. The generated pseudo labels effectively help extend the generalization ability of the detectors.
@@ 3 Grounding DINO
Grounding DINO is a dual-encoder-single-decoder architecture. It contains an image backbone for image feature extraction, a text backbone for text feature extraction, a feature enhancer for image and text feature fusion (Sec. 3.1), a language-guided query selection module for query initialization (Sec. 3.2), and a cross-modality decoder for box refinement (Sec. 3.3). The overall framework is available in Fig. 3.
@@ 3.1 Feature Extraction and Enhancer
Given an (Image, Text) pair, we extract multi-scale image features with an image backbone like Swin Transformer [32], and text features with a text backbone like BERT [8]. Following previous DETR-like detectors [64, 58], multi-scale features are extracted from the outputs of different blocks. After extracting vanilla image and text features, we fed them into a feature enhancer for cross-modality feature fusion. The feature enhancer includes multiple feature enhancer layers. We illustrate a feature enhancer layer in Fig. 3 block 2. We leverage the Deformable self-attention to enhance image features and the vanilla self-attention for text feature enhancers. Inspired by GLIP [26], we add an image-to-text cross-attention and a text-to-image cross-attention for feature fusion. These modules help align features of different modalities.
@@ 3.2 Language-Guided Query Selection
Grounding DINO aims to detect objects from an image specified by an input text. To effectively leverage the input text to guide object detection, we design a language-guided query selection module to select features that are more relevant to the input text as decoder queries. We present the query selection process in Algorithm 1 in PyTorch style. The variables image_features and text_features are used for image and text features, respectively. num_query is the number of queries in the decoder, which is set to $900$ in our implementation. We use bs and ndim for batch size and feature dimension in the pseudo-code. num_img_tokens and num_text_tokens are used for the number of image and text tokens, respectively.
The language-guided query selection module outputs num_query indices. We can extract features based on the selected indices to initialize queries. Following DINO [58], we use mixed query selection to initialize decoder queries. Each decoder query contains two parts: content part and positional part [33], respectively. We formulate the positional part as dynamic anchor boxes [31], which are initialized with encoder outputs. The other part, the content queries, are set to be learnable during training.
@@ 3.3 Cross-Modality Decoder
Algorithm 1 Language-guided query selection.
Input:
image_features:(bs,num_img_tokens,ndim)
text_features:(bs,num_text_tokens,ndim)
num_query:int.
Output:
topk_proposals_idx:(bs,num_query)
logits = torch.einsum("bic,btc->bit",
image_features, text_features)
# bs, num_img_tokens, num_text_tokens
logits_per_img_feat = logits.max(-1)[0]
# bs, num_img_tokens
topk_proposals_idx = torch.topk(
logits_per_image_feature,
num_query, dim=1)[1]
# bs, num_query
@@ 3.5 Loss Function
Grounding-DINO-T (Ours) Swin-T O365 46.7 56.9 / -
Grounding-DINO-T (Ours) Swin-T O365,GoldG 48.1 57.1 / -
Grounding-DINO-T (Ours) Swin-T O365,GoldG,Cap4M 48.4 57.2 / -
Grounding-DINO-L (Ours) Swin-L O365,OI[19],GoldG 52.5 62.6 / 62.7 (63.0 / 63.0)*
Grounding-DINO-L (Ours) Swin-L O365,OI,GoldG,Cap4M,COCO,RefC 60.7 62.6 / -
Table 3: Zero-shot domain transfer to LVIS. *The models are fine-tuned on the LVIS dataset before evaluation.
Model Backbone Pre-Training Data MiniVal [18]
AP APr/APc/APf
MDETR [18]* RN101 GoldG,RefC 24.2 20.9/24.9/24.3
Mask R-CNN [18]* RN101 - 33.3 26.3/34.0/33.9
GLIP-T (C) Swin-T O365,GoldG 24.9 17.7/19.5/31.0
GLIP-T Swin-T O365,GoldG,Cap4M 26.0 20.8/21.4/31.0
Grounding-DINO-T Swin-T O365,GoldG 25.6 14.4/19.6/32.2
Grounding-DINO-T Swin-T O365,GoldG,Cap4M 27.4 18.1/23.3/32.7
Grounding-DINO-L Swin-L O365,OI,GoldG,Cap4M,COCO,RefC 33.9 22.2/30.7/38.8

### v5

@@ Abstract
In this paper, we develop an open-set object detector, called Grounding DINO, by marrying Transformer-based detector DINO with grounded pre-training, which can detect arbitrary objects with human inputs such as category names or referring expressions. The key solution of open-set object detection is introducing language to a closed-set detector for open-set concept generalization. To effectively fuse language and vision modalities, we conceptually divide a closed-set detector into three phases and propose a tight fusion solution, which includes a feature enhancer, a language-guided query selection, and a cross-modality decoder for modalities fusion. We first pre-train Grounding DINO on large-scale datasets, including object detection data, grounding data, and caption data, and evaluate the model on both open-set object detection and referring object detection benchmarks. Grounding DINO performs remarkably well on all three settings, including benchmarks on COCO, LVIS, ODinW, and RefCOCO/+/g. Grounding DINO achieves a $52.5$ AP on the COCO zero-shot [^1] In this paper, ‘zero-shot’ refers to scenarios where the training split of the test dataset is not utilized in the training process. detection benchmark. It sets a new record on the ODinW zero-shot benchmark with a mean $26.1$ AP. We release some checkpoints and inference codes at https://github.com/IDEA-Research/GroundingDINO.
Figure 1: (a) Closed-set object detection requires models to detect objects of pre-defined categories. (b) We evaluate models on novel objects and standard Referring expression comprehension (REC) benchmarks for model generalizations on novel objects with attributes. (c) We present an image editing application by combining Grounding DINO and Stable Diffusion [41]. Best viewed in colors.
@@ 1 Introduction
A key indicator of an Artificial General Intelligence (AGI) system’s capability is its proficiency in handling open-world scenarios. In this paper, we aim to develop a strong system to detect arbitrary objects specified by human language inputs, a task commonly referred to as open-set object detection [^2] We view the terms open-set object detection, open-world object detection, and open-vocabulary object detection the same task in this paper. To avoid confusion, we always use open-set object detection in our paper.. The task has wide applications for its great potential as a generic object detector. For example, we can cooperate with generative models for image editing (as shown in Fig. 1 (b)).
In pursuit of this goal, we design the strong open-set object detector Grounding DINO by following the two principles: tight modality fusion based on DINO [57] and large-scale grounded pre-train for concept generalization.
@@
Tight modality fusion based on DINO. The key to open-set detection is introducing language for unseen object generalization [25, 1, 7]. Most existing open-set detectors are developed by extending closed-set detectors to open-set scenarios with language information. As shown in Fig. 2, a closed-set detector typically has three important modules, a backbone for feature extraction, a neck for feature enhancement, and a head for region refinement (or box prediction). A closed-set detector can be generalized to detect novel objects by learning language-aware region embeddings so that each region can be classified into novel categories in a language-aware semantic space. The key to achieving this goal is using contrastive loss between region outputs and language features at the neck and/or head outputs.
Figure 2: Extending closed-set detectors to open-set scenarios.
To help a model align cross-modality information some work tried to fuse features before the final loss stage. We summarize the modulized design of object detectors in Fig. 2. Feature fusion can be performed in three phases: neck (phase A), query initialization (phase B), and head (phase C). For example, GLIP [25] performs early fusion in the neck module (phase A), and OV-DETR [55] uses language-aware queries as head inputs (phase B). We argue that introducing more feature fusion into the pipeline can facilitate better alignment between different modality features, thereby achieving better performance.
@@
Although conceptually simple, it is hard for previous work to perform feature fusion in all three phases. The design of classical detectors like Faster RCNN makes it hard to interact with language information in most blocks. Unlike classical detectors, the Transformer-based detector method such as DINO has a consistent structure with language blocks. The layer-by-layer design enables it to interact with language information easily. Under this principle, we design three feature fusion approaches in the neck, query initialization, and head phases. More specifically, we design a feature enhancer by stacking self-attention, text-to-image cross-attention, and image-to-text cross-attention as the neck module. We then develop a language-guided query selection method to initialize queries for the detection head. We also design a cross-modality decoder for the head phase with image and text cross-attention layers to boost query representations.
Large-scale grounded pre-train for zero-shot transfer. Most existing open-set models [21, 14] rely on pre-trained CLIP models for concept generalization. Nevertheless, the efficacy of CLIP, specifically pre-trained on image-text pairs, is limited for region-text pair detection tasks, as identified in the RegionCLIP study by RegionCLIP [61]. In contrast, GLIP [25] presents a different way by reformulating object detection as a phrase grounding task and introducing contrastive training between object regions and language phrases on large-scale data. It shows great flexibility for heterogeneous datasets and remarkable performance on closed-set and open-set detection.
We have adopted and refined the grounded training methodology. GLIP’s approach involves concatenating all categories into a sentence in a random order. However, the direct category names concatenation does not consider the potential influence of unrelated categories on each other when extracting features. To mitigate this issue and improve model performance during grounded training, we introduce a technique that utilizes sub-sentence level text features. It removes the attention between unrelated categories during word feature extractions. Further elaboration on this technique can be found in Section 3.4.
We pre-train the Grounding DINO on a large-scale dataset and evaluate the performance on mainstream object detection benchmarks like COCO [29]. While some studies have examined open-set detection models under a "partial label" framework—training on a subset of data (e.g., base categories) and testing on additional categories—we advocate for a fully zero-shot approach to enhance practical applicability. Moreover, we extend the model to another important scenario Referring Expression Comprehension (REC) [34, 30] [^3] We use the term Referring Expression Comprehension (REC) and Referring (Object) Detection exchangeable in this paper., where objects are described with attributes.
@@
We conduct experiments on all three settings, including closed-set detection, open-set detection, and referring object detection, as shown in Fig. 1, to comprehensively evaluate open-set detection performance. Grounding DINO outperforms competitors by a large margin. For example, Grounding DINO reaches a $52.5$ AP on COCO minival without any COCO training data. It also establishes a new state of the art on the ODinW [23] zero-shot benchmark with a $26.1$ mean AP.
@@
Base Detector Fusion (Fig. 2) CLIP Represent. Level (Sec. 3.4) COCO COCO LVIS ODinW RefCOCO/+/g
ViLD [14] Mask R-CNN - sentence partial label partial label
RegionCLIP [61] Faster RCNN - sentence partial label partial label
FindIt [21] Faster RCNN A sentence partial label fine-tune
MDETR [18] DETR A,C word fine-tune zero-shot fine-tune
DQ-DETR [45] DETR A,C word zero-shot fine-tune
GLIP [25] DyHead A word zero-shot zero-shot zero-shot
GLIPv2 [58] DyHead A word zero-shot zero-shot zero-shot
OV-DETR [55] Deformable DETR B sentence partial label partial label
@@
DetCLIP [52] ATSS - sentence zero-shot zero-shot
OmDet [60] Sparse R-CNN C sentence zero-shot
Grounding DINO (Ours) DINO A,B,C sub-sentence zero-shot zero-shot zero-shot zero-shot
Table 1: A comparison of previous open-set object detectors. Our summarization is based on the experiments in their paper, but not the ability to extend their models to other tasks. It is worth noting that some related works may not (only) be designed for the open-set object detection initially, like MDETR [18] and GLIPv2[58], but we list them here for a comprehensive comparison with existing work. We use the term “partial label” for the settings, where models are trained on partial data (e.g. base categories) and evaluated on other cases. [56]
@@ 2 Related Work
Detection Transformers. Grounding DINO is built upon the DETR-like model DINO [57], which is an end-to-end Transformer-based detector. DETR was first proposed in [2] and then has been improved from many directions [64, 33, 13, 5, 48, 17, 4] in the past few years. DAB-DETR [31] introduces anchor boxes as DETR queries for more accurate box prediction. DN-DETR [24] proposes a query-denoising approach to stabilizing the bipartite matching. DINO [57] further develops several techniques including contrastive de-noising and sets a new record on the COCO object detection benchmark. However, such detectors mainly focus on closed-set detection and are difficult to generalize to novel classes because of the limited pre-defined categories.
Open-Set Object Detection. Open-set object detection is trained using existing bounding box annotations and aims at detecting arbitrary classes with the help of language generalization. OV-DETR [56] uses image and text embedding encoded by a CLIP model as queries to decode the category-specified boxes in the DETR framework [2]. ViLD [14] distills knowledge from a CLIP teacher model into a R-CNN-like detector so that the learned region embeddings contain the semantics of language. GLIP [12] formulates object detection as a grounding problem and leverages additional grounding data to help learn aligned semantics at phrase and region levels. It shows that such a formulation can even achieve stronger performance on fully-supervised detection benchmarks. DetCLIP [52] involves large-scale image captioning datasets and uses the generated pseudo labels to expand the knowledge database. The generated pseudo labels effectively help extend the generalization ability.
@@ 3 Grounding DINO
Grounding DINO is a dual-encoder-single-decoder architecture. It contains an image backbone for image feature extraction, a text backbone for text feature extraction, a feature enhancer for image and text feature fusion (Sec. 3.1), a language-guided query selection module for query initialization (Sec. 3.2), and a cross-modality decoder for box refinement (Sec. 3.3).
@@ 3.1 Feature Extraction and Enhancer
Given an (Image, Text) pair, we extract multi-scale image features with an image backbone like Swin Transformer [32], and text features with a text backbone like BERT [8]. Following previous DETR-like detectors [64, 57], multi-scale features are extracted from the outputs of different blocks. After extracting vanilla image and text features, we fed them into a feature enhancer for cross-modality feature fusion. The feature enhancer includes multiple feature enhancer layers. We illustrate a feature enhancer layer in Fig. 3 block 2. We leverage the Deformable self-attention to enhance image features and the vanilla self-attention for text feature enhancers. Inspired by GLIP [25], we add an image-to-text and a text-to-image cross-attention modules for feature fusion. These modules help align features of different modalities.
@@ 3.2 Language-Guided Query Selection
Grounding DINO aims to detect objects from an image specified by an input text. To effectively leverage the input text to guide object detection, we design a language-guided query selection module to select features that are more relevant to the input text as decoder queries.
Let’s denote the image feature as ${{\bf X}}_{I}\in\textsc{R}^{N_{I}\times d}$ and the text features as ${{\bf X}}_{T}\in\textsc{R}^{N_{T}\times d}$. Here, $N_{I}$ represents the number of image tokens, $N_{T}$ indicates the number of text tokens, and $d$ corresponds to the feature dimension. In our experiments, we specifically utilize a feature dimension of $d=256$. Typically, in our models, the value of $N_{I}$ exceeds $10,000$, while $N_{T}$ remains below $256$. Our objective is to extract $N_{q}$ queries from the encoder’s image features to be used as inputs for the decoder. In alignment with the DINO method, we set $N_{q}$ to be $900$. The top $N_{q}$ query indices for the image feature, denoted as ${\bf I}_{N_{q}}$, are selected using the following expression:
${\bf I}_{N_{q}}=\texttt{Top}_{{N_{q}}}(\texttt{Max}^{(-1)}({{\bf X}}_{I}{{\bf X}}_{T}^{\intercal})).$ (1)
In this expression, $\texttt{Top}_{{N_{q}}}$ represents the operation to pick the top $N_{q}$ indices. The function $\texttt{Max}^{(-1)}$ executes the max operation along the 1-1 dimension, and the symbol denotes matrix transposition. We present the query selection process in Algorithm 1 in PyTorch style. The language-guided query selection module outputs $N_{q}$ indices. We can extract features based on the selected indices to initialize queries. Following DINO [57], we use mixed query selection to initialize decoder queries. Each decoder query contains two parts: content part and positional part [33], respectively. We formulate the positional part as dynamic anchor boxes [31], which are initialized with encoder outputs. The other part, the content queries, are set to be learnable during training.
@@ 3.5 Loss Function
Grounding DINO T (Ours) Swin-T O365 46.7 56.9 / -
Grounding DINO T (Ours) Swin-T O365,GoldG 48.1 57.1 / -
Grounding DINO T (Ours) Swin-T O365,GoldG,Cap4M 48.4 57.2 / -
Grounding DINO L (Ours) Swin-L O365,OI[19],GoldG 52.5 62.6 / 62.7 (63.0 / 63.0)*
Grounding DINO L (Ours) Swin-L O365,OI,GoldG,Cap4M,COCO,RefC 60.7 62.6 / -